# MultiverSeg: Scalable Interactive Segmentation of Biomedical Imaging Datasets with In-Context Guidance

### v1

@@ 5 Experimental Setup
We evaluate MultiverSeg and baselines in segmenting a set of images, representing a segmentation task unseen during training. We simulate the process of interactively segmenting each image in a dataset, and adding segmentations to the context set as they are completed.
@@ 5.2 Data
We partition our collection of 78 datasets into 67 datasets for training and 11 datasets for evaluation. We report results on 11 held-out datasets that were unseen by the model during training. These datasets cover 83 tasks and 8 modalities, including unseen image types, anatomies, and labels. The evaluation datasets cover a variety of modalities (MRI, CT, ultrasound, fundus photography, microscopy) and anatomical regions of interest (brain, teeth, bones, abdominal organs, muscles, heart, thorax, cells), including both healthy anatomy and lesions [11, 66, 118, 100, 66, 3, 44, 141, 40, 142, 2].
@@ 5.3 Prompt Simulation
  • Centerline Scribbles: One positive and one negative centerline scribble to start (step 1), followed by one correction centerline scribble per step.
We selected these protocols because center clicks are commonly used for evaluation [114, 130, 76, 47, 77, 59, 10] and centerline scribbles were the most effective scribble type in previous work [126]. We simulate the interactions following ScribblePrompt [126].
@@ 5.4 Baselines
Interactive Segmentation Baselines. We compare to ScribblePrompt [126] a state-of-the-art interactive segmentation method for medical imaging that uses bounding boxes, clicks and scribbles. ScribblePrompt produces more accurate segmentations on unseen biomedical imaging datasets and works with more types of interactions compared to SAM [59] and medical imaging variants [85, 132].
In-Context Segmentation Baselines. We compare to UniverSeg [16], a general state-of-the-art in-context segmentation model that was developed and evaluated on a diverse collection of biomedical images.
Interactive Segmentation in Context Baselines. We construct a new baseline, SP+UVS, by combining UniverSeg [16] and ScribblePrompt [126]. We use the publicly available pre-trained weights for each model. When the context set is empty, we use ScribblePrompt. When provided with a context set, we first make a prediction using UniverSeg, and then refine the prediction with ScribblePrompt.
Consistent with the original published results, we find that UniverSeg has poor performance for small context sets and initializing ScribblePrompt using the UniverSeg prediction hurts performance when the context set is small (Fig. 14). Thus, for context sets with fewer than 5 examples, we ignore the context and only use ScribblePrompt to make predictions.
We did not compare to OnePrompt [128] and LabelAnything [23] because the pre-trained weights were not publicly available and reproducing the training datasets and training pipelines is beyond the scope of this work. Additionally, these models are not designed to perform interactive segmentation in context. We provide more discussion in Sec. C.1.
@@ 5.5 Metrics
We primarily evaluate segmentation quality using Dice score [27], and show 95th percentile Hausdorff distance [49] in Sec. D.2.
Figure 3: Interactions to target Dice on unseen tasks. Number of interactions needed to reach a 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. MultiverSeg requires substantially fewer number of interactions to achieve 90% Dice than the baselines, and as more images are segmented, the average number of interactions required decreases dramatically. Shaded regions show 95% CI from bootstrapping.
Figure 4: Change in total interactions per task. We show average change per task in total number of clicks and scribble steps to segment 18 images to $\geq 90\%$ Dice, comparing MultiverSeg to each baseline method. In all scenarios, MultiverSeg required fewer interactions than both baselines. Error bars show 95% CI from bootstrapping.
Figure 5: Example predictions after 1 interaction step. We show predictions for MultiverSeg and baselines for a random example with a context of 10 examples that were previously segmented to $\geq$90Dice from each held-out task. For each method, we show the prediction after 1 step of interaction: 1 step of centerline scribbles for ACDC [11], BUID [3], and PanDental [1], and 1 center click for SCR [118], WBC [142] and HipXRay [40].
@@ 6 Experiment 1: Evaluation
In this experiment, we evaluate different approaches to segmenting an entire new biomedical dataset. We compare MultiverSeg to ScribblePrompt, which performs interactive segmentation of each image independently, and to SP+UVS, which combines ScribblePrompt with an in-context segmentation model (UniverSeg). We show that MultiverSeg outperforms both techniques.
@@ 6.1 Setup
We evaluate the number of interactions required to achieve a target Dice score on each image using different methods, or a maximum number of interactions if the score was not reached. We use 90% as a target Dice score, because our collection of fully-supervised task-specific nnUNet models achieves an average Dice of $90.15\pm 0.54$ on the same test data. We set 20 center clicks or 10 steps of centerline scribbles as the maximum number of interactions.
For each method and task, we begin by interactively segmenting one randomly sampled image from the training split to $\geq$ 90% Dice using ScribblePrompt [126]. This example is used to seed the context set. We then randomly sample (without replacement) 18 images from the test split, and simulate sequentially segmenting each image.
Data. We report results averaged across 200 simulations for each held-out segmentation task. We exclude tasks with fewer than 18 test examples, leaving 161 tasks from 8 evaluation datasets [2, 11, 3, 1, 40, 118, 142, 125]. We further discuss the choice of this cutoff in Sec. D.1.
@@ 6.2 Results
Interactions per image as a function of dataset size. As more examples are segmented and the context set grows, the number of interactions required to get to $90\%$ Dice (NoI90) on the $\text{n}^{\text{th}}$ example using MultiverSeg decreases substantially (Fig. 3). For ScribblePrompt, NoI90 is approximately constant, because it is not designed to learn from previous examples. With SP+UVS, the number of interactions decreases as more examples are segmented, but it requires more interactions than MultiverSeg. Results by task in Sec. D.2 illustrate a similar trend. Fig. 5 shows predictions for the $10^{th}$ example after 1 step of interaction.
@@
Total interactions. On average, using MultiverSeg reduced the number of scribble steps required to segment each dataset by $(52.70\pm 3.60)\%$ and the number of clicks required by $(35.61\pm 1.95)\%$ (Fig. 4). For larger sets of images, using MultiverSeg results in even greater reductions in the total number of user interactions (Sec. D.2).
@@
Context Set Quality. MultiverSeg was trained with ground truth context set labels. However, at inference time, the context set only includes previously predicted segmentations.
For both MultiverSeg and SP+UVS, thresholding the predictions at 0.5 before adding them to the context set improved the accuracy of predictions for subsequent images. We show results in Sec. D.2 with and without this modification.
@@
As an upper bound on performance, we also evaluated using ground truth labels in the context set instead of predicted segmentations (Sec. D.2). Using ground truth context labels decreases the number of interactions to achieve $90\%$ Dice for both MultiverSeg and SP+UVS, but MultiverSeg still requires fewer interactions.
@@
Bootstrapping In-Context Segmentation. Another approach to segmenting a new dataset is to manually segment an image and then use an in-context segmentation model to segment the rest of the images. We experimented with bootstrapping UniverSeg: starting from a single labeled example as the context set, we sequentially segment each image with UniverSeg and then add it to the context set for the next example. This approach did not produce accurate results ($51.02\pm 0.07$ Dice), likely because UniverSeg has poor performance for small context sets (Fig. 6) and/or context sets with imperfect labels. Because UniverSeg does not have a mechanism to incorporate corrections, it was not possible to achieve 90% Dice for most images. We show results experimenting with this approach in Sec. D.2.2.
@@ 7 Experiment 2: Analysis
When segmenting images sequentially, as in the previous experiment, the performance on the $n^{th}$ image is correlated with the predictions on the previous images. However, in some realistic instances, a few ground truth segmentations might already be available from other previous segmentation efforts. In the following experiments, we analyze MultiverSeg using randomly sampled context sets with ground truth labels. We report results on the test split of 11 evaluation datasets not used during training.
@@ 7.1 In-Context Segmentation
Figure 6: In-context segmentation performance across context set sizes. We compare MultiverSeg to an in-context segmentation method [16] given ground truth context labels. Shading shows 95% CI from bootstrapping.
@@ 8 Conclusion
To enable MultiverSeg, we introduce the first model that can perform interactive segmentation of biomedical images in context. The network segments an image given user interactions and a context set of previously labeled examples. Compared to ScribblePrompt, a state-of-the-art interactive segmentation model, MultiverSeg reduces the number of interactions required to accurately segment a set of images, by as much as 50% on the first 18 images of a dataset.
MultiverSeg opens up new opportunities for research into how best to prioritize sequentially segmenting images from a new dataset. Future research works could improve upon MultiverSeg by investigating better context selection techniques [48, 31, 127] to prioritize labeling images whose labels would be most informative for the segmentation task at hand.
MultiverSeg has the potential to dramatically reduce the manual burden involved in segmenting datasets of biomedical images.
@@ Acknowledgements
Appendix A MultiverSeg Method
A.1 Architecture
@@
Appendix B Data
B.1 Datasets
We build on large dataset gathering efforts like MegaMedical [16, 101, 126] to compile a collection of 78 open-access biomedical imaging datasets for training and evaluation, covering over 54k scans, 16 image types, and 713 labels.
Division of Datasets. The division of datasets and subjects for training, model selection, and evaluation is summarized in Tab. 1. The 78 datasets were divided into 67 training datasets (Tab. 3 and 11 evaluation datasets (Tab. 2). Data from 9 (out of 11) of the evaluation datasets were used for model selection and final evaluation. The other 2 evaluation datasets were completely held-out from model selection and only used in the final evaluation.
Division of Subjects. We split each dataset into 60% train, 20% validation, and 20% test by subject. We used the “train” splits from the 67 training datasets to train MultiverSeg models. We use the “validation” splits from the 67 training datasets and 9 validation datasets to select the best model checkpoint. We report final evaluation results across 11 held-out “test” splits of the 11 evaluation datasets to maximize the diversity of tasks and modalities in our evaluation set (Tab. 2). No data from the 9 validation datasets or 2 test datasets were seen by MultiverSeg during training.
@@ B.1 Datasets
3D Datasets. For 3D modalities, we use the slice with maximum label area (“maxslice”) and the middle slice (“midslice”) for each volume for training of MultiverSeg. For the 3D evaluation datasets (BTCV Cervix [66], ACDC [11], SCD [100], SpineWeb [141], COBRE [2]) we evaluated the slice with the maximum label area for each subject, as in  [126]. We also considered evaluating on the middle slice, as in  [16, 128, 101] and saw similar trends on the validation data. However, we opted for evaluation on maxslices because for our 3D test dataset (COBRE) some labels do not appear in the midslices. Due to the large number of tasks in COBRE, we only consider coronal slices from COBRE for evaluation.
Data Processing. We rescale image intensities to [0,1] and resize images to 128x128.
@@
Table 1: Dataset split overview. Each dataset was split into 60% train, 20% validation and 20% test by subject. Data from the “train” splits of the 67 training datasets were used to train the models. The MultiverSeg models did not see any data from the validation datasets or test datasets during training. Data from the “validation” split of the 9 validation datasets was used for MultiverSeg (MVS) model selection and experimenting with different evaluation methods of baselines. We report final results on the held-out test splits of 11 evaluation datasets: data from the “test” splits of the 9 validation datasets and the “test” splits of the 2 test datasets. To train the fully-supervised nnUNet baselines, we used the training and validation splits of the 11 evaluation datasets.
@@
Test Datasets 2 nnUNet training nnUNet training Final evaluation
Table 2: Evaluation datasets. We assembled the following set of datasets to evaluate MultiverSeg and baseline methods. For the relative size of datasets, we include the number of unique scans (subject and modality pairs) and labels that each dataset has. These datasets were unseen by MultiverSeg during training. Three datasets were completely held-out from model selection. The validation splits of the other 9 datasets were used for selecting the best model checkpoint. We report final results on the test splits of these 11 datasets.
@@
B.2 Synthetic Task Generation

### v2

@@ 5 Experimental Setup
We evaluate MultiverSeg and baselines in segmenting a set of images, representing a segmentation task unseen during training. We simulate the process of interactively segmenting each image in a dataset, and of adding the segmentations to the context set as they are completed.
@@ 5.2 Data
We partition our collection of 79 datasets into 67 datasets for training and 12 datasets held-out for evaluation. We report results on the 12 held-out datasets that were unseen by the model during training. These datasets cover 187 tasks and 8 modalities, including unseen image types, anatomies, and labels. The evaluation datasets cover a variety of modalities (MRI, CT, ultrasound, fundus photography, microscopy) and anatomical regions of interest (brain, teeth, bones, abdominal organs, muscles, heart, thorax, cells), including both healthy anatomy and lesions [11, 68, 122, 103, 68, 3, 46, 145, 42, 146, 2, 129].
@@ 5.3 Prompt Simulation
  • Centerline Scribbles: One positive and one negative centerline scribble to start (step 1), followed by one positive or negative correction centerline scribble per step.
We selected these protocols because center clicks are commonly used for evaluation [118, 134, 78, 49, 79, 61, 10] and centerline scribbles were the most effective prompt in [130].
@@ 5.4 Baselines
Interactive Segmentation Baselines. We compare to five interactive segmentation methods trained on biomedical images: ScribblePrompt [130], MedSAM [88], SAM-Med2D [19], and IMIS-Net [20]. We also evaluated two general interactive segmentation methods, SAM [61] and SegNext [79], which were trained on natural images. We focus on ScribblePrompt over SAM, and medical imaging variants of SAM [88, 19, 20], because it produced more accurate segmentations on unseen biomedical imaging datasets [130, 108] and has faster inference runtime.
In-Context Segmentation Baselines. We compare to UniverSeg [16], a general state-of-the-art in-context segmentation model that was trained on a diverse collection of biomedical images. We did not compare to OnePrompt [132] and LabelAnything [25] because the pre-trained weights were not publicly available. We discuss further in Sec. D.1.
Interactive In-Context Segmentation Baselines. We construct a new baseline, SP+UVS, by combining UniverSeg [16] and ScribblePrompt [130]. We use the publicly available pre-trained weights for each model. When the context set is empty, we use ScribblePrompt. When provided with a context set, we first predict using UniverSeg, and then refine the prediction with ScribblePrompt.
Consistent with the original published results, we find that UniverSeg has poor performance for small context sets and initializing ScribblePrompt using the UniverSeg prediction hurts performance when the context set is small (Fig. 16). Thus, for context sets with fewer than 5 examples, we ignore the context and use only ScribblePrompt to make predictions.
@@ 5.5 Metrics
We evaluate segmentation quality using Dice score [29], and show 95th percentile Hausdorff distance [51] in Sec. E.2. MultiverSeg, UniverSeg, and ScribblePrompt were all trained and developed on images at $128^{2}$ resolution. Unless otherwise noted, we evaluated on images resized to $256^{2}$ resolution to demonstrate performance at a higher, more realistic, resolution. We show results with similar trends in Sec. E.5, evaluating at $128^{2}$ resolution.
Figure 3: Interactions to target Dice on unseen tasks. Number of interactions needed to reach a 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. MultiverSeg requires substantially fewer number of interactions to achieve 90% Dice than the baselines, and as more images are segmented, the average number of interactions required decreases dramatically.
Figure 4: Interactions per image by unseen dataset. We show average number of clicks and scribble steps per image to segment 18 images to $\geq 90\%$ Dice for each method. In all scenarios, MultiverSeg required fewer or the same number of interactions than the best baseline. Error bars show 95% CI from bootstrapping.
Figure 5: Example predictions after 1 interaction step. We show predictions for MultiverSeg and the top two performing baselines on a randomly chosen example from each held-out task. We use a context set of 10 examples that were previously segmented to $\geq$90% Dice. For each method, we show the prediction after 1 step of interaction: 1 step of centerline scribbles for ACDC [11], BUID [3], and PanDental [1], and 1 center click for SCR [122], WBC [146], HipXRay [42], and TotalSegmentator [129].
@@ 6 Experiment 1: Evaluation
In this experiment, we evaluate different approaches to segmenting an entire new biomedical dataset. We compare MultiverSeg to ScribblePrompt, SAM, SegNext, SAM-Med2D, IMIS-Net, and MedSAM, which perform interactive segmentation of each image independently, and to SP+UVS, which combines ScribblePrompt with an in-context segmentation model (UniverSeg). We show that MultiverSeg outperforms all of the baselines.
@@ 6.1 Setup
We evaluate the number of interactions required to achieve a target Dice score on each image using different methods, or a maximum number of interactions if the score was not reached. We use 90% as a target Dice score, because our collection of fully-supervised task-specific nnUNet models achieves an average Dice of $88.67\pm 0.47$ on the same test data. We set 20 center clicks or 10 steps of centerline scribbles as the maximum number of interactions.
For each method and task, we begin by interactively segmenting one randomly sampled image from the training split to $\geq$ 90% Dice using ScribblePrompt. This example is used to seed the context set. We then randomly sample (without replacement) 18 images from the test split, and simulate sequentially segmenting each image.
Data. We report results averaged across 200 simulations for each held-out segmentation task. We exclude tasks with fewer than 18 test examples, leaving 161 tasks from 8 evaluation datasets [2, 11, 3, 1, 42, 122, 146, 129]. We further discuss the choice of this cutoff in Sec. E.1.
@@ 6.2 Results
Interactions per image as a function of dataset size. As more examples are segmented and the context set grows, the number of interactions required to get to $90\%$ Dice (NoI90) on the $\text{n}^{\text{th}}$ example using MultiverSeg decreases substantially (Fig. 3). For interactive segmentation methods, NoI90 is approximately constant, because they are not designed to learn from previous examples. With SP+UVS, the number of interactions decreases as more examples are segmented, but it requires more interactions than MultiverSeg. Results by task in Sec. E.2 show a similar trend. Fig. 5 shows predictions for the $10^{th}$ example after 1 step of interaction.
@@
Total interactions. On average, using MultiverSeg reduced the number of clicks required to segment each dataset by $(36.41\pm 1.33)\%$ and the number of scribbles steps required by $(25.26\pm 1.80)\%$ compared to ScribblePrompt (Fig. 4). For larger sets of images, using MultiverSeg results in even greater reductions in the total number of user interactions (Sec. E.2).
Other Baselines. MedSAM, which only works with bounding boxes, had an average Dice of $65.93\pm 4.82$ and was only able to reach 90% Dice for $5.6\%$ of examples. SegNext failed with scribbles due to GPU memory limits.
@@
Context Set Quality. MultiverSeg was trained with ground truth context set labels. However, at inference time, the context set only includes previously predicted segmentations. For both MultiverSeg and SP+UVS, thresholding the predictions at 0.5 before adding them to the context set improved the accuracy of predictions for subsequent images. We show the effect of this modification in Sec. E.2.
@@
As an upper bound on performance, we also evaluated using ground truth labels in the context set instead of predicted segmentations (Sec. E.2). Using ground truth context labels decreases the number of interactions to achieve $90\%$ Dice for both MultiverSeg and SP+UVS, but MultiverSeg still requires fewer interactions.
@@
Bootstrapping In-Context Segmentation. Another approach to segmenting a new dataset is to manually segment an image, and then use an in-context segmentation model to segment the rest of the images. We experimented with bootstrapping UniverSeg: starting from a single labeled example as the context set, we sequentially segment each image with UniverSeg and then add it to the context set for the next example. This approach did not produce accurate results ($48.89\pm 1.87$ Dice), likely because UniverSeg has poor performance for small context sets (Fig. 6) and/or context sets with imperfect labels. Because UniverSeg does not have a mechanism to incorporate corrections, it was not possible to achieve 90% Dice for most images. We show results experimenting with this approach in Sec. E.3.
Task-Specific Fine-Tuning. Another approach is to interactively segment a few images, and then fine-tune ScribblePrompt using those labeled examples to produce a task-specific interactive segmentation model. This requires computational overhead and machine-learning expertise that is often unavailable in biomedical research or clinical workflows. As we show in Sec. E.4, even if it were practical, the fine-tuned models do not perform as well as MultiverSeg. Fine-tuning each task-specific model took 20 minutes on average using a NVIDIA A100 GPU. In contrast, MultiverSeg’s inference runtime is <0.15<0.15 seconds, even with a context set size of 64 examples (Sec. F.3).
Limitations. MultiverSeg does not perform as well on tasks where the context set images vary substantially in composition, especially for limited set sizes. E.g., Fig. 11 shows that with clicks, MultiverSeg underperforms ScribblePrompt on the BUID dataset until the context set has $\geq 5$ examples. Given scribbles, which provide more information, context is less helpful, and ScribblePrompt and MultiverSeg have similar performance on BUID (Fig. 12).
@@ 7 Experiment 2: Analysis
When segmenting images sequentially, as in the previous experiment, the performance on the $n^{th}$ image is correlated with the predictions on the previous images. However, in some realistic instances, a few ground-truth segmentations might be available from other previous segmentation efforts. In the following experiments, we analyze MultiverSeg using randomly sampled context sets with ground truth labels. We report results on the test split of 12 evaluation datasets not used during training. Because of the computational burden of training 187 task-specific nnUNets, we trained on images with $128^{2}$ resolution. Thus, for this experiment, we report results at $128^{2}$ resolution.
@@ 7.1 In-Context Segmentation
Figure 6: In-context segmentation performance across context set sizes. We compare MultiverSeg to an in-context segmentation method, UniverSeg [16], given ground truth context labels. Shading shows 95% CI from bootstrapping.
@@ 8 Conclusion
To enable MultiverSeg, we introduce the first model that can perform interactive segmentation of biomedical images in context. The network segments an image given user interactions and a context set of previously labeled examples. Compared to ScribblePrompt, a state-of-the-art interactive segmentation model, MultiverSeg reduces the number of clicks required to accurately segment a set of images, by 36% on average on the first 18 images of a dataset.
MultiverSeg opens up new opportunities for research into how best to prioritize sequentially segmenting images from a new dataset. Future research works could improve upon MultiverSeg by investigating better context selection techniques [50, 33, 131] to prioritize labeling images whose labels would be most informative for the segmentation task at hand. MultiverSeg has the potential to dramatically reduce the manual burden involved in segmenting datasets of biomedical images.
@@ Acknowledgements
Appendix A Code
Code and pre-trained weights are available at https://multiverseg.csail.mit.edu.
Appendix B MultiverSeg Method
B.1 Architecture
@@
Appendix C Data
C.1 Datasets
We build on large dataset gathering efforts like MegaMedical [16, 104, 130] to compile a collection of 79 open-access biomedical imaging datasets for training and evaluation, covering over 54k scans, 16 image types, and 713 labels.
Division of Datasets. The division of datasets and subjects for training, model selection, and evaluation is summarized in Tab. 1. The 79 datasets were divided into 67 training datasets (Tab. 3 and 12 evaluation datasets (Tab. 2). Data from 9 (out of 12) of the evaluation datasets were used for model selection and final evaluation. The other 3 evaluation datasets were completely held-out from model selection and only used in the final evaluation.
Division of Subjects. We split each dataset into 60% train, 20% validation, and 20% test by subject. We used the “train” splits from the 67 training datasets to train MultiverSeg models. We use the “validation” splits from the 67 training datasets and 9 validation datasets to select the best model checkpoint. We report final evaluation results across 12 held-out “test” splits of the 12 evaluation datasets to maximize the diversity of tasks and modalities in our evaluation set (Tab. 2). No data from the 9 validation datasets or 3 test datasets were seen by MultiverSeg during training.
@@ C.1 Datasets
3D Datasets. For 3D modalities, we use the slice with maximum label area (“maxslice”) and the middle slice (“midslice”) for each volume for training of MultiverSeg. For the 3D evaluation datasets (BTCV Cervix [68], ACDC [11], SCD [103], SpineWeb [145], COBRE [2], TotalSegmentator [129]) we evaluated the slice with the maximum label area for each subject, as in  [130]. We also considered evaluating on the middle slice, as in  [16, 132, 104] and saw similar trends on the validation data. However, we opted for evaluation on maxslices because for our 3D test datasets (COBRE, TotalSegmentator) some labels do not appear in the midslices. Due to the large number of tasks in COBRE and TotalSegmentator, we only consider coronal slices from these datasets for evaluation.
Data Processing and Image Resolution. We rescale image intensities to [0,1], padded square with zeros. For training, we resized images to $128^{2}$. In our final evaluations (Sec. 6), we use images resized to $256^{2}$. We show additional evaluations on $128^{2}$ sized images in Sec. E.5.
@@
Table 1: Dataset split overview. Each dataset was split into 60% train, 20% validation and 20% test by subject. Data from the “train” splits of the 67 training datasets were used to train the models. The MultiverSeg models did not see any data from the validation datasets or test datasets during training. Data from the “validation” split of the 9 validation datasets was used for MultiverSeg (MVS) model selection and experimenting with different evaluation methods of baselines. We report final results on the held-out test splits of 12 evaluation datasets: data from the “test” splits of the 9 validation datasets and the “test” splits of the 2 test datasets. To train the fully-supervised nnUNet baselines, we used the training and validation splits of the 12 evaluation datasets.
@@
Test Datasets 3 nnUNet training nnUNet training Final evaluation
Table 2: Evaluation datasets. We assembled the following set of datasets to evaluate MultiverSeg and baseline methods. For the relative size of datasets, we include the number of unique scans (subject and modality pairs) and labels that each dataset has. These datasets were unseen by MultiverSeg during training. Three datasets were completely held-out from model selection. The validation splits of the other 9 datasets were used for selecting the best model checkpoint. We report final results on the test splits of these 12 datasets.
@@
TotalSegmentator [129] 104 anatomic structures (27 organs, 59 bones, 10 muscles, and 8 vessels) 1,204 104 CT
@@
C.2 Synthetic Task Generation